# MTIL: Encoding Full History with Mamba for Temporal Imitation Learning

### v1

@@ Abstract
Standard imitation learning (IL) methods have achieved considerable success in robotics, yet often rely on the Markov assumption, limiting their applicability to tasks where historical context is crucial for disambiguating current observations. This limitation hinders performance in long-horizon sequential manipulation tasks where the correct action depends on past events not fully captured by the current state. To address this fundamental challenge, we introduce Mamba Temporal Imitation Learning (MTIL), a novel approach that leverages the recurrent state dynamics inherent in State Space Models (SSMs), specifically the Mamba architecture. MTIL encodes the entire trajectory history into a compressed hidden state, conditioning action predictions on this comprehensive temporal context alongside current multi-modal observations. Through extensive experiments on simulated benchmarks (ALOHA, Robomimic, LIBERO) and on challenging real-world sequential manipulation tasks that requiring long-term temporal reasoning, MTIL significantly outperforms state-of-the-art methods like ACT and Diffusion Policy. Our findings affirm the necessity of full temporal context for robust sequential decision-making and validate MTIL as a powerful approach that transcends the inherent limitations of Markovian imitation learning. Our implementation is available at MTIL
Keywords: Imitation Learning, Robot Manipulation, Sequence Modeling, State Space Models
1 Introduction
Imitation Learning (IL) has emerged as a powerful paradigm for teaching robots complex skills directly from expert demonstrations, bypassing the need for intricate reward engineering often required in reinforcement learning [1, 2, 3]. Behavioral Cloning (BC), the simplest form of IL, learns a direct mapping from observations to actions via supervised learning and has enabled robots to perform a variety of tasks [2, 3, 4, 1]. Recent advancements, particularly leveraging powerful sequence models and generative approaches, have led to state-of-the-art (SOTA) methods such as the Action Chunking Transformer (ACT) [5, 6] and Diffusion Policy [7, 8, 9, 10], which excel at learning visuomotor control policies for complex manipulation.
@@ 1 Introduction
Despite these successes, a fundamental limitation persists in many current IL approaches: the reliance on the Markov assumption. These methods typically predict the action $a_{t}$ based solely on the current observation $o_{t}$ or a short, fixed-length history window $o_{t-k:t}$  [1, 11, 12, 13]. This assumption breaks down in tasks where the history beyond this limited window is necessary to resolve ambiguity in the current observation. Consider a sequential task requiring a robot to first place an object at location A, and subsequently move it to location B. At an intermediate configuration, the robot’s visual observation and proprioceptive state ($o_{t}$) might be identical regardless of whether it has successfully completed the sub-task at location A. A Markovian policy, lacking the memory of visiting A, cannot distinguish these fundamentally different historical contexts and may erroneously proceed directly to B, failing the task [1, 14, 15]. This temporal ambiguity signifies an underlying Partially Observable Markov Decision Process (POMDP), posing a critical challenge for standard IL methods in state-dependent tasks.
@@
Since ambiguous tasks manifest as POMDPs and human demonstrations are inherently non-Markovian, effective imitation necessitates history-aware policies.While the critical role of historical context has been increasingly recognized  [12], effectively and efficiently encoding comprehensive trajectory histories into actionable representations remains a significant challenge. Addressing this, we introduce Mamba Temporal Imitation Learning (MTIL).This approach is specifically designed to incorporate the complete observational history into decision-making by leveraging the unique properties of State Space Models (SSMs), particularly the recently developed Mamba architecture [16, 17]. Mamba’s recurrent structure allows it to maintain a compressed hidden state $h_{t}$ that theoretically encapsulates information from the entire preceding observation sequence $H_{t}=(o_{1},...,o_{t})$. Instead of relying solely on $o_{t}$, MTIL learns a policy $\pi(a_{t}|h_{t},o_{t})$ that explicitly conditions the action prediction on this history-infused hidden state $h_{t}$ in conjunction with the current observation $o_{t}$, enabling differentiation between observationally similar states and the correct execution of complex sequential tasks. Our contributions are threefold:
@@
    A novel Mamba-based IL architecture, MTIL, is proposed, specifically designed for full trajectory history encoding via its recurrent state to overcome limitations inherent in Markovian sequential task learning.
@@
    The fundamental challenge of temporal ambiguity in IL is addressed by incorporating comprehensive historical context, offering a promising pathway towards learning and executing long-horizon, state-dependent manipulation sequences previously intractable for methods reliant on limited context.
@@
    Extensive empirical validation is provided across diverse benchmarks (simulation and real-world), demonstrating significant performance improvements over existing state-of-the-art methods (including ACT, Diffusion Policy), particularly on tasks demanding temporal reasoning and long-term memory.
2 Related Works
2.1 Markovian and Short-History Imitation Learning
@@
2.2 Temporal and Sequential Imitation Learning
The inadequacy of the strict Markov assumption has long motivated efforts to incorporate temporal context. Early explorations employed Recurrent Neural Networks (RNNs) like LSTMs [1, 4, 12, 24]. However, these architectures face challenges with long-term dependencies due to vanishing gradients [24]. Furthermore, practical implementations often resorted to fixed history windows and periodic state resets (e.g., sequence lengths of 10-50 in Robomimic [12, 25]), precluding the capture of full trajectory history. More recently, Transformer-based models (e.g., BeT [26], RT-1 [27, 28], OPTIMUS [29], ICRT [30], Baku [27, 28], MDT [31]) have become prominent, utilizing self-attention to model sequence correlations. Yet, the $O(L^{2})$ computational complexity of attention imposes practical limits on the size of context window  [1, 24], hindering their ability to efficiently process and encode entire long trajectories in the manner afforded by linear-complexity SSMs. Distinct strategies for managing long horizons involve temporal abstraction. Hierarchical Imitation Learning [14, 15, 32, 33, 34] and Skill Chaining [14, 35, 36] decompose tasks, learning policies over skills or sub-goals. Waypoint-based methods like AWE [19] or primitive-based approaches like PRIME [15] operate at higher levels of abstraction. While effective, these approaches fundamentally differ from MTIL, which aims to directly model the complete low-level observation-action sequence history, potentially offering robustness against issues like error propagation in skill chaining [14, 36].
2.3 State Space Models (SSMs) and Mamba in Robotics
@@
3 Mamba Temporal Imitation Learning (MTIL)
@@
Figure 1: Overview of the Mamba Temporal Imitation Learning (MTIL) architecture. Multi-modal inputs (images via DINOv2, state) are fused and processed by sequential Mamba-2 blocks, updating the recurrent state $h_{t}$ which encodes history. At each step $t$ across the entire trajectory, MTIL predicts an action chunk $\hat{a}_{t:t+K-1}$ (current plus $K-1$ future steps). This is supervised via L2 loss against ground truth actions $a_{t:t+K-1}$ from the demonstration (using last action for padding when near trajectory end). The historical context embedded in $h_{t}$ enables temporally coherent, long-horizon action generation.
3.1 Background and Motivation
@@
Directly computing or representing the belief state $b_{t}$ is generally intractable. This motivates learning a compressed history representation $h_{t}\approx b_{t}$ using recurrent models. State Space Models (SSMs) like Mamba [16] offer a particularly compelling approach, providing a structured recurrent update $h_{t}=f(h_{t-1},x_{t})$ (where $x_{t}$ encodes $o_{t}$) with linear time complexity $O(L)$. This efficiency is critical for tractably encoding the long sequences required for full history representation, overcoming limitations of prior architectures [47].
Our motivation for MTIL stems from leveraging Mamba’s state $h_{t}$ as this potent, efficiently computed representation of the full history. By conditioning actions on both the current observation $o_{t}$ and this history-infused state $h_{t}$, MTIL learns a non-Markovian policy:
@@
3.2 Leveraging Full Trajectory History with Mamba-2
@@
Crucially, the input-dependent parameters $(\Delta_{t},\bar{\mathbf{B}}_{t},\mathbf{C}_{t})=f(x_{t})$ enable selective state dynamics, allowing the model to dynamically modulate information flow and preserve salient historical context. This selective mechanism, combined with inherent linear-time complexity $O(L)$, facilitates learning from complete trajectory histories—a significant advantage over quadratic-complexity $O(L^{2})$ attention mechanisms or memory-limited RNNs for long sequences. The resulting state $h_{t}$ acts as a dynamic summary of the salient history $(x_{1},...,x_{t-1})$, furnishing the requisite context even when the instantaneous observation $x_{t}$ is ambiguous. The MTIL policy leverages this directly:
@@
0:  Expert trajectories $\mathcal{D}=\{\tau_{i}\}$, $\tau_{i}=(o_{1},a_{1},...,o_{T_{i}},a_{T_{i}})$, MTIL Policy $\pi_{\theta}$, initialized parameters $\theta$, Loss $\mathcal{L}$ (MSE), Chunk size $k$, Optimizer $Opt$
@@
3.3 MTIL Training and Inference
@@
MTIL enables imitation learning across complete expert trajectories, utilizing the architecture outlined in Figure 1. Distinctively, MTIL employs a sequential training procedure (Algorithm 1). This step-wise paradigm, leveraging Mamba’s recurrent ‘step‘ function, is fundamental to efficiently encoding arbitrarily long trajectory histories from high-dimensional observations (e.g., images) within feasible memory constraints—a key departure from parallel window-based approaches.At each timestep $t$ along the trajectory (length $T$), the policy receives the observation embedding $x_{t}$, updates its history-encoding hidden state from $h_{t-1}$ to $h_{t}$, and predicts a chunk of $K$ actions $\hat{a}_{t:t+K-1}$, covering the current and $K-1$ future steps. Learning proceeds by minimizing the Mean Squared Error (MSE) between this predicted chunk $\hat{a}_{t:t+K-1}$ and the ground truth actions $a_{t:t+K-1}$ sourced from the demonstration data. Notably, when $t+K-1$ exceeds the trajectory length $T$, the ground truth chunk is padded by repeating the final demonstrated action $a_{T}$. During inference (Algorithm 2), the trained policy operates autoregressively. It uses the same ‘step‘ function to update its hidden state $h_{t}$ based on the current observation and predict the action chunk $\hat{a}_{t:t+K-1}$. For enhanced stability and smoothness in execution, temporal aggregation strategies [19, 20, 5] are commonly applied. This involves computing the final action for time $t$ by applying a weighted average over multiple predictions made for this specific timestep but generated at different preceding steps.
4 Experimental Results
We conducted extensive experiments to evaluate the performance of MTIL across various benchmarks and real-world scenarios.See more details in Appendix A
4.1 ACT benchmark
Table 1: Success rates(%) on ACT.
Cube Bimanual
Method Transfer Insertion
ACT [5] 90 50
MTIL (10-step) 92 56
MTIL (Full) 100 84
MTIL’s efficiency and performance were evaluated on ACT [5] simulation tasks. Table 1 presents the final success rates, while Figure 2 shows learning curves and a backbone ablation. Compared to ACT [5], MTIL leveraging full trajectory history (approx.400 steps in these tasks) achieves substantially higher success rates (Table 1: 100%/84% vs.90%/50% for ACT on Cube Transfer/Bimanual Insertion) and markedly faster convergence (Figure 2a). We also evaluated MTIL variants using shorter fixed history windows (10, 20, and 50 steps),and their final performance did not show statistically significant differences from each other (10-step results: 92%/56%), confirming that such short-history modeling is clearly inferior to utilizing the full-history available. Furthermore, to isolate the benefits of the full-history approach from visual feature quality, we conducted an ablation comparing MTIL using DINOv2 [49] against ResNet18 [50] (which is used by ACT). While DINOv2 offers stronger features—a factor contributing to its selection, particularly for real-world robustness—the results (Figure 2b) combined with the short-versus-full history comparison underscore that the primary performance leap stems from MTIL’s core mechanism: encoding comprehensive trajectory context.
@@ 4.1 ACT benchmark
4.2 LIBERO Benchmark

### v3

@@ Abstract
Standard imitation learning (IL) methods have achieved considerable success in robotics, yet often rely on the Markov assumption, which falters in long-horizon tasks where history is crucial for resolving perceptual ambiguity. This limitation stems not only from a conceptual gap but also from a fundamental computational barrier: prevailing architectures like Transformers are often constrained by quadratic complexity, rendering the processing of long, high-dimensional observation sequences infeasible. To overcome this dual challenge, we introduce Mamba Temporal Imitation Learning (MTIL). Our approach represents a new paradigm for robotic learning, which we frame as a practical synthesis of World Model and Dynamical System concepts. By leveraging the linear-time recurrent dynamics of State Space Models (SSMs), MTIL learns an implicit, action-oriented world model that efficiently encodes the entire trajectory history into a compressed, evolving state. This allows the policy to be conditioned on a comprehensive temporal context, transcending the confines of Markovian approaches. Through extensive experiments on simulated benchmarks (ACT, Robomimic, LIBERO) and on challenging real-world tasks, MTIL demonstrates superior performance against SOTA methods like ACT and Diffusion Policy, particularly in resolving long-term temporal ambiguities. Our findings not only affirm the necessity of full temporal context but also validate MTIL as a powerful and a computationally feasible approach for learning long-horizon, non-Markovian behaviors from high-dimensional observations. Project code are available at https://github.com/yulinzhouZYL/MTIL
I Introduction
@@
Imitation Learning (IL) has emerged as a powerful paradigm for teaching robots complex skills directly from expert demonstrations, bypassing the need for intricate reward engineering often required in reinforcement learning [1, 2, 3]. Behavioral Cloning (BC), the simplest form of IL, learns a direct mapping from observations to actions via supervised learning and has enabled robots to perform a variety of tasks [2, 3, 4, 1]. Recent advancements, particularly leveraging powerful sequence models and generative approaches, have led to state-of-the-art (SOTA) methods such as the Action Chunking Transformer (ACT) [5, 6] and Diffusion Policy [7, 8, 9, 10], which excel at learning visuomotor control policies for complex manipulation. Despite these successes, a fundamental limitation persists in many current IL approaches: the reliance on the Markov assumption. These methods typically predict the action $a_{t}$ based solely on the current observation $o_{t}$ or a short, fixed-length history window $o_{t-k:t}$  [1, 11, 12, 13]. This assumption breaks down in tasks where the history beyond this limited window is necessary to resolve ambiguity in the current observation. Consider a sequential task requiring a robot to first place an object at location A, and subsequently move it to location B. At an intermediate configuration, the robot’s visual observation and proprioceptive state ($o_{t}$) might be identical regardless of whether it has successfully completed the sub-task at location A. A Markovian policy, lacking the memory of visiting A, cannot distinguish these fundamentally different historical contexts and may erroneously proceed directly to B, failing the task [1, 14, 15]. This temporal ambiguity signifies an underlying Partially Observable Markov Decision Process (POMDP), posing a critical challenge for standard IL methods in state-dependent tasks.
@@
Since ambiguous tasks manifest as POMDPs and human demonstrations are inherently non-Markovian, effective imitation necessitates history-aware policies. While the critical role of historical context has been increasingly recognized [12], a fundamental barrier has remained: the computational infeasibility of processing long, high-dimensional observation histories with prevailing architectures like Transformers. Addressing this, we introduce Mamba Temporal Imitation Learning (MTIL). Our approach is specifically designed to incorporate the complete observational history into decision-making by leveraging the unique properties of State Space Models (SSMs), particularly the recently developed Mamba architecture [16, 17]. Mamba’s recurrent structure allows it to maintain a compressed hidden state $h_{t}$ that theoretically encapsulates information from the entire preceding observation sequence $H_{t}=(o_{1},...,o_{t})$. Instead of relying solely on $o_{t}$, MTIL learns a policy $\pi(a_{t}|h_{t},o_{t})$ that explicitly conditions the action prediction on this history-infused hidden state $h_{t}$ in conjunction with the current observation $o_{t}$, enabling differentiation between observationally similar states and the correct execution of complex sequential tasks. Our contributions are threefold:
@@
    We propose MTIL, a novel architecture that is the first to leverage the linear-time recurrence of State Space Models to make full-trajectory imitation learning from high-dimensional visual data computationally feasible on commodity hardware, breaking the quadratic bottleneck of attention-based models.
@@
    We provide a new theoretical framing for this approach, positing MTIL as learning an implicit dynamical system. This system’s evolving state acts as a continuous belief-state representation, offering a robust solution to the core problem of temporal ambiguity in partially observable environments.
@@
    We provide extensive empirical validation demonstrating that MTIL significantly outperforms state-of-the-art methods, including ACT and Diffusion Policy. Furthermore, on tasks requiring long-term memory, MTIL also surpasses other full-history-capable baselines like Transformer-XL, validating the unique advantages of its underlying architecture.
II Related Works
II-A Markovian and Short-History Imitation Learning
@@
II-B Temporal and Sequential Imitation Learning
The inadequacy of the strict Markov assumption has long motivated efforts to incorporate temporal context. Early explorations employed Recurrent Neural Networks (RNNs) like LSTMs [1, 4, 12, 24]. However, these architectures face challenges with long-term dependencies due to vanishing gradients [24]. Furthermore, practical implementations often resorted to fixed history windows and periodic state resets (e.g., sequence lengths of 10-50 in Robomimic [12, 25]), precluding the capture of full trajectory history. More recently, Transformer-based models (e.g., BeT [26], RT-1 [27, 28], OPTIMUS [29], ICRT [30], Baku [27, 28], MDT [31]) have become prominent, utilizing self-attention to model sequence correlations. However, the $O(L^{2})$ computational complexity of attention imposes practical limits on the size of the context window [1, 24], hindering their ability to efficiently process entire long trajectories. Even recurrent variants like Transformer-XL [32], while theoretically capable of processing long sequences, still rely on the computationally intensive attention mechanism. Distinct strategies for managing long horizons involve temporal abstraction. Hierarchical Imitation Learning [14, 15, 33, 34, 35] and Skill Chaining [14, 36, 37] decompose tasks, learning policies over skills or sub-goals. Waypoint-based methods like AWE [19] or primitive-based approaches like PRIME [15] operate at higher levels of abstraction. While effective, these approaches fundamentally differ from MTIL, which aims to directly model the complete low-level observation-action sequence history, potentially offering robustness against issues like error propagation in skill chaining [14, 37].
II-C State Space Models (SSMs) and Mamba in Robotics
@@
III Mamba Temporal Imitation Learning (MTIL)
@@
Figure 1: Overview of the Mamba Temporal Imitation Learning (MTIL) architecture. Multi-modal inputs (images via DINOv2, state) are fused and processed by sequential Mamba-2 blocks, updating the recurrent state $h_{t}$ which encodes history. At each step $t$ across the entire trajectory, MTIL predicts an action chunk $\hat{a}_{t:t+K-1}$ (current plus K1K-1 future steps). This is supervised via L2 loss against ground truth actions $a_{t:t+K-1}$ from the demonstration (using last action for padding when near trajectory end). The historical context embedded in $h_{t}$ enables temporally coherent, long-horizon action generation.
III-A Background and Motivation
@@
Directly computing or representing the belief state $b_{t}$ is generally intractable. This motivates learning a compressed history representation $h_{t}\approx b_{t}$ using recurrent models. This aligns with the core ideas of both World Models, which learn a predictive latent state of the world, and Dynamical Systems (DS) approaches to control, which rely on an evolving internal state. State Space Models (SSMs) like Mamba [16] offer a particularly compelling synthesis of these ideas. They provide a structured recurrent update $h_{t}=f(h_{t-1},x_{t})$ (where $x_{t}$ encodes $o_{t}$) with linear time complexity $O(L)$. This efficiency is the critical enabler for tractably encoding the long sequences required for full history representation, overcoming the computational barriers of prior architectures [48].
Our motivation for MTIL stems from leveraging Mamba’s state $h_{t}$ as this potent, efficiently computed representation of the full history. We view $h_{t}$ as the state of a learned, implicit dynamical system that acts as an action-oriented world model. By conditioning actions on both the current observation $o_{t}$ and this history-infused state $h_{t}$, MTIL learns a non-Markovian policy:
@@
III-B Leveraging Full Trajectory History with Mamba-2
@@
Crucially, the input-dependent parameters $(\Delta_{t},\bar{\mathbf{B}}_{t},\mathbf{C}_{t})=f(x_{t})$ enable selective state dynamics. This allows the model to learn a highly non-linear dynamical system where the system matrices themselves adapt based on the current input. This selective mechanism, combined with inherent linear-time complexity $O(L)$, facilitates learning from complete trajectory histories—a significant advantage over quadratic-complexity $O(L^{2})$ attention mechanisms. The resulting state $h_{t}$ acts as a dynamic summary of the salient history $(x_{1},...,x_{t-1})$, furnishing the requisite context even when the instantaneous observation $x_{t}$ is ambiguous. The MTIL policy leverages this directly:
@@
0:  Expert trajectories $\mathcal{D}=\{\tau_{i}\}$, $\tau_{i}=(o_{1},a_{1},...,o_{T_{i}},a_{T_{i}})$, MTIL Policy $\pi_{\theta}$, initialized parameters $\theta$, Loss $\mathcal{L}$ (MSE), Chunk size $k$, Optimizer OptOpt
@@
III-C MTIL Training and Inference
@@
MTIL enables imitation learning across complete expert trajectories, utilizing the architecture outlined in Figure 1. Distinctively, MTIL employs a sequential training procedure (Algorithm 1). This step-wise paradigm, leveraging Mamba’s recurrent ‘step‘ function, is fundamental to efficiently encoding arbitrarily long trajectories from high-dimensional observations (e.g., images) within feasible memory constraints—a key departure from parallel window-based approaches.A naive implementation of this sequential process would be limited to a batch size of one, posing a challenge for training efficiency. To address this, we introduce a novel batch-parallel training scheme. Instead of processing a single trajectory, our method processes a batch of $B$ trajectories simultaneously. At each timestep $t$, the model takes a batch of observations, updates their respective hidden states and computes the loss concurrently. This approach preserves the crucial temporal integrity within each trajectory while fully leveraging the parallel processing power of modern GPUs, making MTIL’s training time competitive with highly-parallelizable Markovian methods.During training, at each timestep $t$, the policy receives the observation embedding $x_{t}$, updates its history-encoding state from $h_{t-1}$ to $h_{t}$, and predicts an action chunk $\hat{a}_{t:t+K-1}$. Learning proceeds by minimizing the Mean Squared Error (MSE) against the ground truth actions $a_{t:t+K-1}$. During inference (Algorithm 2), the trained policy operates autoregressively, using the same ‘step‘ function to update its state and predict actions. For enhanced stability, temporal aggregation strategies [19, 20, 5] are applied, averaging over predictions from overlapping action chunks to produce a smoother final action.
IV Experimental Results
We conducted extensive experiments to evaluate the performance of MTIL across various benchmarks and real-world scenarios.All results stem from a rigorous protocol over three random seeds (100, 200, 300), with 50 roll-outs and the checkpoint for each run selected based on the minimum validation loss or as the final success rate for Robomimic.
IV-A ACT benchmark
TABLE I: Success Rates (%) on the ACT Benchmark. Results are averaged over 3 seeds, All experiments run on a single RTX 4090.
Method History Length Cube Transfer (%) Bimanual Insertion (%)
ACT [5] 1 (Markovian) 90.0 $\pm$ 2.0 50.0 $\pm$ 3.5
Diffusion Policy [7] 1 (Markovian) 72.0 $\pm$ 2.6 28.0 $\pm$ 3.2
Diffusion Policy [7] 10 78.0 $\pm$ 2.5 32.0 $\pm$ 4.1
Diffusion Policy [7] 20 80.0 $\pm$ 2.2 34.0 $\pm$ 3.8
Diffusion Policy [7] 30 82.0 $\pm$ 2.1 36.0 $\pm$ 3.5
Diffusion Policy [7] 40 OOM
Transformer-XL [32] Full (400) 86.0 $\pm$ 2.5 42.0 $\pm$ 4.0
MTIL (10-step) 10 92.0 $\pm$ 1.5 56.0 $\pm$ 2.5
MTIL (Full) Full (400) 100.0 $\pm$ 0.0 84.0 $\pm$ 2.1
@@ IV-A ACT benchmark
We evaluated MTIL on the ACT benchmark to dissect its performance, efficiency, and learning dynamics against prominent architectural paradigms. The results, which juxtapose success rates with architectural choices and history lengths, are presented in Table I. The findings decisively establish MTIL’s superiority. On both tasks, MTIL (Full) achieves a perfect or near-perfect success rate, drastically outperforming all baselines. The learning curves in Figure 2(a) illuminate this outcome, showing that MTIL not only attains a higher performance ceiling but also converges significantly faster, indicating a more stable and sample-efficient learning process. Conversely, the performance of attention-based models reveals a critical insight: naively incorporating history is an inefficient, and ultimately, a computationally infeasible strategy. While Diffusion Policy’s success rate scales with history length, it remains notably inferior to the simple Markovian ACT baseline and incurs a prohibitive computational cost, culminating in an Out-of-Memory (OOM) error. Even Transformer-XL, theoretically capable of full-history processing, fails to match ACT, reinforcing the hypothesis that attention is a suboptimal inductive bias for modeling the continuous dynamics of physical interaction. Furthermore, the backbone ablation in Figure 2(b) confirms our advantage is architectural. MTIL, even with an identical ResNet18 backbone [50], substantially outperforms ACT. The use of a stronger DINOv2 backbone [51] further widens this gap.This proves MTIL’s success stems from a fundamentally superior paradigm: a computationally efficient recurrent architecture that is intrinsically better suited to capturing the temporal fabric of the physical world.
IV-B LIBERO Benchmark